# Dealing with Logs and Zeros in Regression Models

### v2

@@ 1 Introduction
Our estimator, iterated OLS (iOLS), is the theoretically justified analogue of the ad-hoc transformation $\log(1+Y)$ transformation commonly used to accommodate zeros. Whereas this popular fix imposes an arbitrary constant shift, iOLS replaces this shift with a data-driven adjustment that is updated iteratively through OLS until convergence. This procedure yields an estimator that is nearly as simple to implement as the popular fix but is consistent for the structural parameter of the exponential mean model. More generally, we show that iOLS solves the same objective function as Gamma Pseudo-Maximum Likelihood (GPML), via a sequence of fixed-point updates. Our key methodological contribution is to reinterpret the exponential regression model as a simple iterative linear regression model, thereby making available the full machinery of linear models, such as partialling out fixed effects and using instrumental variables. [^1] Although our framework focuses on GPML, it naturally extends to any power variance function (Bar-Lev and Enis, 1986), i.e. the class of Tweedie distributions with a log link including PPML and the compound Poisson-Gamma. A systematic study of their properties is left for future research.
@@
To document the ubiquity of log specifications, the presence of zeros, and the lack of consensus on how to handle them, we collected evidence from three sources: (i) a review of empirical publications, (ii) a survey conducted in seminars, and (iii) an online research forum. First, we reviewed all articles published in the American Economic Review (AER) between 2016 and 2020. Figure 1 summarizes our findings. Log-linear and log-log models rank among the most frequent specifications: nearly 40% of empirical papers employed a log specification, and 36% of these faced the problem of the log of zero. Several solutions were used in practice. Most authors retained zero observations and either (1) added a positive discretionary constant (48%), (2) employed Poisson Pseudo Maximum Likelihood (PPML) or GPML (35%), or (3) applied the inverse hyperbolic sine (15%). Discarding non-positive observations occurred in 31% of publications, and in about 20% of cases authors compared multiple methods to assess robustness. [^2] This excludes cases where a linear specification was used as a default. See Table D.2 in the Appendix for details on data collection. Second, we conducted a survey during three online seminars in economics departments, asking researchers: “What would you do when facing the log of zero?” The distribution of responses closely mirrored the AER evidence, except for a stronger stated preference for mixture models. [^3] Appendix D provides details about the survey design and responses. Third, the issue extends well beyond economics. A ResearchGate thread titled “Log transformation of values that include 0 (zero) for statistical analyses?”, posted in 2014, had been viewed over 120,000 times and had attracted 38 contributions from researchers in medicine, biology, statistics, and engineering by August 2020. The suggested solutions were again broadly consistent with those observed in economics.
@@ 2.3 Iterated OLS
Let us make the following standard assumptions about $Y_{i}$, $X_{i}$ and the error term $U_{i}$.
@@
For any fixed $\delta<\infty$, we can simply choose $c_{i}(\beta,\delta)=E[\log(U_{i}+\delta)]$,101010 In practice, one must use an empirical analog $\hat{c}(\delta,\beta)$ obtained as follows. Let us write $X_{i}^{\prime}\beta=\alpha+X_{i1}^{\prime}\beta_{1}$, with $\beta=(\alpha,\beta_{1}^{\prime})^{\prime}$ where $\alpha$ is the constant term and $X_{i1}^{\prime}\beta_{1}$ represents the non-deterministic part. From $E[U_{i}]=1$ it is easy to obtain $\alpha(\beta_{1})=\log(E[Y_{i}\exp(-X_{i1}^{\prime}\beta_{1})])$ then to characterize the desired term as a function of $\beta_{1}$: $c(\delta,\beta)=E[\log(Y_{i}+\delta\exp(\alpha(\beta_{1})+X_{i1}^{\prime}\beta_{1}))-\alpha(\beta_{1})-X_{i1}^{\prime}\beta_{1}]$, finally replace by the empirical counterparts. so that under Assumption 2, we have $E[\overline{\upsilon}_{i}]=0$ and $E[\overline{\upsilon}_{i}|X_{i}]=o(\delta^{-2})$. The solution of the fixed-point equation $\hat{\beta}=X^{\prime}X^{-1}X^{\prime}\tilde{Y}(\hat{\beta},\delta)$ is thus a biased estimator of $\beta$, denoted $\beta_{\delta}=\beta+o(\delta^{-2})$.
@@
To overcome this difficulty, we adopt a two-phase strategy: the first phase solves a sequence of approximate problems for increasing values of $\delta$, where the log transformation ensures global and fast convergence to increasingly accurate approximations. The second phase then applies the limiting transformation in (16), starting from the Phase-1 solution, to eliminate the residual $o(\delta^{-2})$ bias. This design ensures both global convergence and consistency.
@@
Under Assumptions 1 and 2, the iOLS estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, iOLS is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega)$, as $n\to\infty$, where the covariance can be consistently estimated using the asymptotic covariance of the OLS estimator in the last iteration up to minor modifications.
This theorem guarantees $\sqrt{n}$-consistent estimates and, for any fixed $n$, each iterative process converges after a finite number of iterations: $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping. The numerical convergence will hence be slower for larger sample sizes $n$ and modulus $\kappa$ closer to 1, where $\kappa$ depends on the DGP and is affected by the choice of the sequence $\delta$ and stability parameter $\rho$.
@@ 2.4 Endogeneity: Iterated 2SLS
Under Assumptions 3, 4, the i2SLS estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}^{IV}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, the i2SLS estimator is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}^{IV}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega^{IV})$, as $n\to\infty$, where the asymptotic covariance can be consistently estimated using the asymptotic covariance of the 2SLS estimator in the last iteration up to minor modifications.
@@
2.5 Partialling-out fixed-effects
Nonlinear regression models with multi-way fixed effects face two main challenges: the large number of nuisance parameters and the IPP. We address both by avoiding direct estimation of the fixed effects. Specifically, we apply a Frisch-Waugh-Lovell projection to partial out all fixed-effect indicators, as detailed in the next algorithm.
Algorithm 3 (Fixed-effects)
We introduce two minor modifications to Algorithm 1. First, let us partition the regressor matrix as $X=[X_{0},X_{1}]$, where $X_{0}$ comprises all fixed-effect dummies and $X_{1}$ contains the remaining covariates (including the intercept). At initialization, we project $X_{1}$ onto the orthogonal complement of $X_{0}$ once and for all. Thereafter, the algorithm proceeds exactly as before, except that at each iteration we also partial out $\tilde{Y}$ from $X_{0}$. This approach also applies to Algorithm 2.
@@ 2.5 Partialling-out fixed-effects
where $M_{0}$ denotes the projection onto the orthogonal complement of $\text{span}(X_{0})$,
In this formulation, the transformed error $M_{0}\overline{\upsilon}_{i}$ satisfies the necessary exogeneity condition of Proposition 1, so that the large-sample properties established in Theorem 1 carry over with only minor modifications. Importantly, because the projection on $X_{0}$ is computed once and reused throughout, the presence of high-dimensional fixed effects has essentially no impact on the computational complexity of the iterative procedure.
Proposition 1
Under Assumption 2, and $E[M_{0}X_{1}|X_{0}]=0$, the transformed error term $M_{0}\overline{\upsilon}_{i}$ satisfies the exogeneity restriction $E[(M_{0}X_{1})^{\prime}M_{0}\overline{\upsilon}]=o(\delta^{-2})$.
@@
The additional orthogonality condition $E[M_{0}X_{1}\mid X_{0}]=0$ ensures that, once additive fixed effects are partialled out from $X_{1}$, the residual variation in $X_{1}$ is uncorrelated with interactions of those same fixed effects. This condition is automatically satisfied in a one-way fixed-effects model, since $X_{0}^{\prime}\beta_{0}$ is constant within each cell. With multi-way fixed effects, however, the nonlinear transformation $\exp(X_{0}^{\prime}\beta_{0})$ implicitly generates cross-products of fixed effects that vary within cell and cannot be captured by the linear projection. For instance, if unit and time dummies both enter the model, then a within-transformation of the error will itself contain products of unit and time effects. The resulting projection error is hence proportional to the within-cell variance of $\exp(X_{0}^{\prime}\beta_{0})$. A Taylor expansion shows that the linear component of $X_{0}^{\prime}\beta_{0}$ is absorbed by the projection, leaving residual terms driven by curvature. Thus, when $\exp(X_{0}^{\prime}\beta_{0})$ spans only a narrow range of values within cell, the bias is negligible.
One way to avoid this potential bias is to estimate one dimension of the fixed effects directly (e.g., time effects) while applying the orthogonal projection only to the other dimension (e.g., unit effects). In that setup, the transformed error no longer contains cross-terms between units and periods and therefore retains the same exogeneity properties as in the one-way case. [^12] This strategy does not induce an IPP provided that the dimension with few observations is purged. For instance, in a panel with $T=2$ and $n=10$, one should estimate the time dummies and remove the unit dummies, which would otherwise be poorly estimated given only two time periods.

### v3

@@ 1 Introduction
Our estimator, iterated OLS (iOLS), is the theoretically justified analogue of the ad-hoc $\log(1+Y)$ transformation commonly used to accommodate zeros. Whereas this popular fix imposes an arbitrary constant shift, iOLS replaces this shift with a data-driven adjustment that is updated iteratively through OLS until convergence. This procedure yields an estimator that is nearly as simple to implement as the popular fix but is consistent for the structural parameter of the exponential mean model. More generally, we show that iOLS solves the same objective function as Gamma Pseudo-Maximum Likelihood (GPML), via a sequence of fixed-point updates. Our key methodological contribution is to reinterpret the exponential regression model as a simple iterative linear regression model, thereby making available the full machinery of linear models, such as partialling out fixed effects and using instrumental variables. [^1] Our framework naturally extends to any power variance function (Bar-Lev and Enis, 1986) beyond GPML, i.e. the class of Tweedie distributions with a log link including PPML and the compound Poisson-Gamma. A systematic study of their properties is left for future research.
@@
To document the ubiquity of log specifications, the presence of zeros, and the lack of consensus on how to handle them, we collected evidence from three sources: (i) a review of empirical publications, (ii) a survey conducted in seminars, and (iii) an online research forum. First, we reviewed all articles published in the American Economic Review (AER) between 2016 and 2020. Figure 1 summarizes our findings. Log-linear and log-log models rank among the most frequent specifications: nearly 40% of empirical papers employed a log specification, and 36% of these faced the problem of the log of zero. Several solutions were used in practice. Most authors retained zero observations and either (1) added a positive discretionary constant (48%), (2) employed Poisson Pseudo Maximum Likelihood (PPML) or GPML (35%), or (3) applied the inverse hyperbolic sine (15%). Discarding non-positive observations occurred in 31% of publications, and in about 20% of cases authors compared multiple methods to assess robustness. [^2] This excludes cases where a linear specification was used as a default. See Table E.2 in the Appendix for details on data collection. Second, we conducted a survey during three online seminars in economics departments, asking researchers: “What would you do when facing the log of zero?” The distribution of responses closely mirrored the AER evidence, except for a stronger stated preference for mixture models. [^3] Appendix E provides details about the survey design and responses. Third, the issue extends well beyond economics. A ResearchGate thread titled “Log transformation of values that include 0 (zero) for statistical analyses?”, posted in 2014, had been viewed over 120,000 times and had attracted 38 contributions from researchers in medicine, biology, statistics, and engineering by August 2020. The suggested solutions were again broadly consistent with those observed in economics.
@@ 2.3 Iterated OLS
Let us make the following assumptions about $Y_{i}$, $X_{i}$ and the error term $U_{i}$.
@@
Assumption 3 (Overlap)
Let $D_{Y}:=\operatorname{diag}\big(\mathbf{1}\{Y_{1}>0\},\ldots,\mathbf{1}\{Y_{n}>0\}\big)$. There exists $\alpha\in(0,1]$ such that, for every nonzero $v\in\mathbb{R}^{K}$, $v^{\prime}X^{\prime}D_{Y}Xv\ \geq\ \alpha\ v^{\prime}X^{\prime}Xv$. Equivalently, letting $s_{i}(v):=v^{\prime}X_{i}$, $\frac{\sum_{i:\,Y_{i}>0}s_{i}(v)^{2}}{\sum_{i=1}^{n}s_{i}(v)^{2}}\ \geq\ \alpha.$
Assumption 3 requires that, in every regressor direction $v$, a fixed fraction $\alpha$ of the sample variability $v^{\prime}X_{i}$ is contributed by observations with $Y_{i}>0$. This rules out boundary patterns where regressors vary only when $Y_{i}=0$. In particular, for a binary regressor $X\in\{0,1\}$, the condition excludes the degenerate case where $X_{i}=1$ only when $Y_{i}=0$.
@@
For any fixed $\delta<\infty$, we can simply choose $c_{i}(\beta,\delta)=E[\log(U_{i}+\delta)]$, [^11] In practice, one must use an empirical analog $\hat{c}(\delta,\beta)$ obtained as follows. Let us write $X_{i}^{\prime}\beta=\alpha+X_{i1}^{\prime}\beta_{1}$, with $\beta=(\alpha,\beta_{1}^{\prime})^{\prime}$ where $\alpha$ is the constant term and $X_{i1}^{\prime}\beta_{1}$ represents the non-deterministic part. From $E[U_{i}]=1$ it is easy to obtain $\alpha(\beta_{1})=\log(E[Y_{i}\exp(-X_{i1}^{\prime}\beta_{1})])$ then to characterize the desired term as a function of $\beta_{1}$: $c(\delta,\beta)=E[\log(Y_{i}+\delta\exp(\alpha(\beta_{1})+X_{i1}^{\prime}\beta_{1}))-\alpha(\beta_{1})-X_{i1}^{\prime}\beta_{1}]$, finally replace by the empirical counterparts. so that under Assumption 2, we have $E[\overline{\upsilon}_{i}]=0$ and $E[\overline{\upsilon}_{i}|X_{i}]=O(\delta^{-2})$. The solution of the fixed-point equation $\hat{\beta}=X^{\prime}X^{-1}X^{\prime}\tilde{Y}(\hat{\beta},\delta)$ is thus a biased estimator of $\beta$, denoted $\beta_{\delta}=\beta+O(\delta^{-2})$.
@@
To overcome this difficulty, we adopt a two-phase strategy: the first phase solves a sequence of approximate problems for increasing values of $\delta$, where the log transformation ensures global and fast convergence to increasingly accurate approximations. The second phase then applies the limiting transformation in (16), starting from the Phase-1 solution, to eliminate the residual $O(\delta^{-2})$ bias. This design ensures both uniform global convergence and consistency.
@@
Under Assumptions 1, 2 and 3, the iOLS estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, iOLS is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega)$, as $n\to\infty$, where the covariance can be consistently estimated using the asymptotic covariance of the OLS estimator in the last iteration up to minor modifications.
This theorem guarantees $\sqrt{n}$-consistent estimates and, for any fixed $n$, each iterative process converges after a finite number of iterations: $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping. The numerical convergence will hence be slower for larger sample sizes $n$ and modulus $\kappa$ closer to 1, where $\kappa$ depends on the DGP and is affected by the choice of the sequence for $\delta$ and stability parameter $\rho$.
@@ 2.4 Endogeneity: Iterated 2SLS
Assumption 6 (Overlap)
There exists $\alpha\in(0,1]$ such that, for every nonzero $v\in\mathbb{R}^{K}$, $v^{\prime}X^{\prime}P_{Z}D_{Y}P_{Z}Xv\ \geq\ \alpha\;v^{\prime}X^{\prime}P_{Z}Xv$.
Assumption 6 requires that, in every instrumented regressor direction $v$, a fixed fraction $\alpha$ of the sample variability is contributed by observations with $Y_{i}>0$. Again, this rules out separation patterns where the instrumented part of $X$ varies only when $Y_{i}=0$.
@@
Under Assumptions 4, 5, and 6, the i2SLS estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}^{IV}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, the i2SLS estimator is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}^{IV}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega^{IV})$, as $n\to\infty$, where the asymptotic covariance can be consistently estimated using the asymptotic covariance of the 2SLS estimator in the last iteration up to minor modifications.
@@
2.5 Fixed-effects
Nonlinear regressions with multi-way fixed effects (FE) face two related challenges: a high-dimensional set of nuisance parameters and an IPP when some FE dimensions are short. We tackle both using Frisch-Waugh-Lovell projections to partial out FE indicators. We develop three complementary algorithms tailored to the empirical design: (i) when every FE dimension is large, iOLS-HDFE handles the high dimensionality but would suffer an IPP if any FE dimension were short; (ii) when FE dimensions are short, iOLS-FD removes them ex ante and iterates entirely in the residualized space, thereby avoiding the IPP; and (iii) with a mix of large and short FE dimensions, iOLS-FD-HDFE estimates the large FE dimensions and partials out the short ones.
2.5.1 All fixed-effects dimensions are large
First, let us partition the regressor matrix as $X=[X_{0},X_{1}]$, where $X_{0}$ comprises all fixed-effect dummies and $X_{1}$ contains the remaining covariates (including the intercept). Let us denote the aggregate fixed-effect term by $\Lambda=X_{0}^{\prime}\beta_{0}$, and the linear projections $P_{0}$ (onto $span(X_{0})$) and its orthogonal complement $M_{0}$.
Algorithm 3 (iOLS-HDFE)
[…]
  1. 1.
    Initialize $t=0$ and let $\hat{\beta}_{0}$ and $\hat{\Lambda}_{0}$ be initial estimates;
  2. 2.
    Transform $Y$ into $\tilde{Y}(\hat{\beta}_{t})=\log(Y+\delta\exp(X_{1}^{\prime}\hat{\beta}_{t}+\hat{\Lambda}_{t}))-c(\delta,\hat{\beta}_{t},\hat{\Lambda}_{t})$;
  3. 3.
    Partial-out $\tilde{Y}(\hat{\beta}_{t},\hat{\Lambda}_{t})$ into $\acute{\tilde{Y}}=M_{0}\tilde{Y}(\hat{\beta}_{t},\hat{\Lambda}_{t})$ and $X_{1}$ into $\acute{X_{1}}=M_{0}X_{1}$;
  4. 4.
    Run OLS: $\hat{\beta}_{t+1}=[\acute{X_{1}}^{\prime}\acute{X_{1}}]^{-1}\acute{X_{1}}^{\prime}\acute{\tilde{Y}}$;
  5. 5.
    Recover: $\hat{\Lambda}_{t+1}=P_{0}\left(\tilde{Y}-X_{1}^{\prime}\hat{\beta}_{t+1}\right)$, and update $t$ to $t+1$;
  6. 6.
    Iterate steps (2.) to (5.) until $\hat{\beta}_{t}$ converges. Call the result $\hat{\beta}_{\delta}$.
  7. 7.
    Re-initialize $t=0$, update $\delta(i)$ to $\delta(i+1)$ and iterate steps (2.) to (6.) until $\hat{\beta}_{\delta(i)}$ converges. Call the solution $\hat{\beta}_{\delta}$. Proceed to Phase 2 to eliminate the remaining approximation bias $O(\delta^{-2})$ as in Algorithm 1.
This estimator is consistent and asymptotically normal only if every fixed-effect (FE) dimension grow sufficiently fast with the sample size, and some modified overlap condition. If any FE dimension remains fixed, an incidental-parameter problem (IPP) emerges. [^13] A formal proof of this estimator’s asymptotic properties parallels that of Algorithm 5 and is hence omitted. The issue can be observed from the estimating equations.
The transformed regression used to update $\beta$ is
$M_{0}\,\log\!\Big(\tfrac{Y+\delta\exp(X_{1}^{\prime}\beta+\Lambda)}{\exp(c(\beta,\delta))}\Big)\;=\;M_{0}X_{1}^{\prime}\beta\;+\;M_{0}\overline{\upsilon},$ (17)
while the fixed effects are updated using the projection onto $span(X_{0})$:
$\Lambda\;=\;P_{0}\,\log\!\Big(\tfrac{Y+\delta\exp(X_{1}^{\prime}\beta+\Lambda)}{\exp(c(\beta,\delta))}\Big)\;-\;P_{0}X_{1}^{\prime}\beta-P_{0}\Lambda\;-\;P_{0}\overline{\upsilon},$ (18)
with $P_{0}\Lambda=0$. The term $P_{0}\overline{\upsilon}$ is the sample group means of the transformed error. Under Assumption 2, these group means converge to zero whenever each FE group size grows to infinity with the sample, so (18) identifies $\Lambda$ asymptotically and the noise carried into the $\beta$-update through $\hat{\Lambda}$ disappears. However, if at least one FE dimension is short, the associated group means do not converge to zero, and neither does $P_{0}\overline{\upsilon}$, leaving asymptotic bias in $\hat{\Lambda}$. This noise contaminates the $\beta$-update and produces an IPP: even as $n\to\infty$, the limiting distribution of $\hat{\beta}$ is biased. This is the same mechanism behind the IPP in PPML: with a short FE dimension, the projection step cannot “average out” the FE component.
2.5.2 All fixed-effects dimensions are short
We now introduce an estimator that removes the short fixed-effects dimensions ex-ante and thereby sidesteps the IPP. The idea is simple: apply the projection $M_{0}$ to both the covariates and the transformed outcome, and then iterate OLS entirely in the residualized space. No fixed effect is ever estimated, so noise from short FE groups cannot feed back into $\hat{\beta}$. Theorem 1 applies with minor modifications, assuming the following overlap condition.
Assumption 7 (Residualized overlap)
Let $D_{Y}:=\operatorname{diag}\big(\mathbf{1}\{Y_{1}>0\},\ldots,\mathbf{1}\{Y_{n}>0\}\big)$. There exists $\alpha_{C}\in(0,1]$ such that, for every nonzero $v\in\mathbb{R}^{p}$, $v^{\prime}\,\acute{X}_{1}^{\prime}D_{Y}\acute{X}_{1}\,v\ \geq\ \alpha_{C}\;v^{\prime}\,\acute{X}_{1}^{\prime}\acute{X}_{1}\,v.$
Assumption 7 requires that a non-trivial share of the within-group sample variation of the residualized regressors is carried by units with $Y_{i}>0$. This rules out within-group separation patterns.
Algorithm 4 (iOLS-FD)
[…]
  1. 1.
    Initialize $t=0$ and let $\hat{\beta}_{0}$ be initial estimate;
  2. 2.
    Partial-out $X_{1}$ into $\acute{X_{1}}=M_{0}X_{1}$;
  3. 3.
    Transform $Y$ into $\acute{\tilde{Y}}(\hat{\beta}_{t})=M_{0}\left(\log(Y+\delta\exp(\acute{X_{1}}\hat{\beta}_{t}))-c(\delta,\hat{\beta}_{t})\right)$;
  4. 4.
    Run OLS: $\hat{\beta}_{t+1}=[\acute{X_{1}}^{\prime}\acute{X_{1}}]^{-1}\acute{X_{1}}^{\prime}\acute{\tilde{Y}}$, and update $t$ to $t+1$;
  5. 5.
    Iterate steps (3.) and (4.) until $\hat{\beta}_{t}$ converges. Call the result $\hat{\beta}_{\delta}$.
  6. 6.
    Re-initialize $t=0$, update $\delta(i)$ to $\delta(i+1)$ and iterate steps (2.) to (6.) until $\hat{\beta}_{\delta(i)}$ converges. Call the solution $\hat{\beta}_{\delta}$. Proceed to Phase 2 to eliminate the remaining approximation bias $O(\delta^{-2})$ as in Algorithm 1.
@@ 2.5.2 All fixed-effects dimensions are short
where the projector $M_{0}$ is computed once and reused throughout, so the presence of high-dimensional fixed effects has very limited impact on the computational complexity of the iterative procedure. The asymptotic properties are established in Theorem 3.
Theorem 3 (Consistency and Normality of iOLS-FD)
Under Assumptions 1, 2, 7, and $E[M_{0}X_{1}|X_{0}]=0$, the transformed error term $M_{0}\overline{\upsilon}$ satisfies the necessary exogeneity restriction $E[(M_{0}X_{1})_{i}(M_{0}\overline{\upsilon})_{i}]=O(\delta^{-2})$. Moreover, the iOLS-FD estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, iOLS-FD is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega)$, as $n\to\infty$, where the covariance can be consistently estimated using the asymptotic covariance of the OLS estimator in the last iteration up to minor modifications.
@@
The properties of iOLS established in Theorem 1 hence carry over to iOLS-FD, under the added assumption $E[M_{0}X_{1}\mid X_{0}]=0$. This additional orthogonality condition ensures that, once additive fixed effects are partialled out from $X_{1}$, the residual variation in $X_{1}$ is uncorrelated with interactions of those same fixed effects. This condition is automatically satisfied in a one-way fixed-effects model, since $X_{0}^{\prime}\beta_{0}$ is constant within each cell. With multi-way fixed effects, however, the nonlinear transformation $\exp(X_{0}^{\prime}\beta_{0})$ implicitly generates cross-products of fixed effects that vary within cell and cannot be captured by the linear projection. For instance, if unit and time dummies both enter the model, then a within-transformation of the error will itself contain products of unit and time effects. The resulting projection error is hence proportional to the within-cell variance of $\exp(X_{0}^{\prime}\beta_{0})$. This bias is hence negligible only if $\exp(X_{0}^{\prime}\beta_{0})$ spans a narrow range of values within cell.